a. The prices of yen are quoted in Tokyo at ¥115/$ Bid and ¥118/$ Ask. What are the direct quotations of the bid and ask prices in New York?  

b. If the exchange rates are quoted as ¥145/€ and HK$20/€, then what is the cross- exchange rate of yen per HK$ (¥/HK$)?
Answer #1

Answer a)

Direct Quotes in Tokyo = 1$ = 115Yen - 118Yen

Direct Quote in Neew York = 1 Yen = 1/118$ - 1/115$

1 Yen = 0.00847 $ - 0.00870$

Answer b)

1 Euro = 145 Yen (yen / Euro)

1 Euro = HK$20 (HK$ / Euro) Indirect will be 1HK$ = 1/20 Euro OR 0.05 Euro (Euro / HK$)

Yen / HKS = Yen / Euro * Euro/ HK$

substituting the respective values

Yen / HKS = 145 * 0.05

= 7.25

1 HK$ = 7.25 Yen